# FlowBotHD: History-Aware Diffuser Handling Ambiguities in Articulated Objects Manipulation

### v1

@@ 5.3 Articulation Manipulation Policy
Along with the history-aware diffusion model, we propose a novel articulation manipulation policy that applies the flow predicting model to real manipulation tasks taking into consideration the multi-modality of diffusion and the incorporation of object history.
@@
Switch Grasp Point (SG): Following Eisner et al. [1], to achieve the best manipulation efficiency with a suction gripper, we want to grasp at the point with the largest leverage; thus the optimal action given the predicted flow is to grasp the point that has the maximum flow norm [1]. Due to the multi-modality of diffusion, the prediction at different steps will vary; each prediction might imply a different point that the robot should optimally grasp. In order to correct for past potential grasping mistakes as well as to maintain a reasonable re-grasping cost, we propose to switch the grasp point whenever the leverage - defined as the current largest flow norm - of the newly predicted grasp point is larger than that of the current grasp point by a threshold: $||\hat{F}||_{max}-||\hat{F}||_{g}>\delta l$, where $\hat{F}$ is the current step’s flow prediction for all points in the point cloud, $g$ is the index of the point that is closest to the current grasp point, and $\delta l$ is the threshold which we set to 0.2 in practice.
@@
Consistency Check (CC): Being able to model multi-modality, some samples from the diffusion model will likely be correct, whereas other samples will be incorrect. With the expectation that diffusion will produce the correct prediction given enough samples, we design a heuristic consistency check to filter the actions that are not consistent with the previously executed motion. Specifically, we define the consistency direction $g_{his}$ as the gripper movement of the previous execution (we only keep the previous execution if its motion was greater than 0). To perform a consistent following step, the predicted action’s direction shouldn’t deviate too much from that of the previous step; thus we set an angle threshold $\delta\theta_{cc}$ for the consistency check: $\arccos{\frac{\hat{f}\cdot g_{his}}{||\hat{f}||\cdot||g_{his}||}}<\delta\theta_{cc}$, where $\hat{f}$ is the current sampled flow direction. We will sample actions from the diffusion model until we find one that passes the consistency check, before executing the sampled action.
@@
History Filter (HF): Since we have trained our model to use the history flow prediction to guide the current step’s predictions, we don’t want previous mistakes to mislead the model. Therefore, we include a history filter that only updates the history point cloud $\textbf{P}_{his}$ and history flow $F_{his}$ if the previous action achieves a good movement in execution. Specifically, after an execution, we only update the history if two criteria are satisfied:
@@
    the delta gripper motion is consistent with past executions: $\arccos{\frac{\delta g_{cur}\cdot g_{his}}{||\delta g_{cur}||\cdot||g_{his}||}}<90^{\circ}$ where $g_{cur}$ is the current gripper movement and $g_{his}$ is the gripper movement from the previous execution.
@@
    the gripper motion is larger than a threshold $||\delta g_{cur}||>\delta m$ where $g_{cur}$ is the current gripper movement and $\delta m$ is the threshold
@@
     $\textbf{P}_{obs}\leftarrow$ Observation
     $\hat{F}\sim D_{\theta}(\textbf{P}_{obs},\textbf{S}_{his})$   Sample the current flow direction from the diffusion model
     $\hat{f}\leftarrow\texttt{GetMaxFlow}(\hat{F})$
@@
     $p_{grasp}\leftarrow\texttt{SelectGraspPoint}(\hat{F})$
@@
        $p_{grasp,his}=p_{grasp}$, ;
@@
        $\texttt{Grasp}(P_{grasp})$ ;
@@
     $g_{cur}\leftarrow\texttt{GetDeltaGripper}()$
     if $\texttt{Consistent}(g_{cur},g_{his})$  then
        $g_{his}\leftarrow g_{cur}$
        if $||g_{cur}||>\delta m$ then
           $\textbf{S}_{his}\leftarrow(\textbf{P}_{obs},\hat{F})$ ;
@@ 6.1 Simulation results
To evaluate each model’s ability to open objects, we build a simulation environment with PyBullet [42] that simulates operating a suction gripper to operate objects modeled in PartNet-Mobility dataset. We initialize every object as fully closed and filtered out instances that don’t open from the fully closed state even with the ground truth action sequence. Following prior work [1], we compute 2 evaluation metrics: Success Rate: the percent of objects the model succeeds to open $>90\%$ of the maximum opening amount, and Normalized Distance [13]: the mean distance between the end state of this trial and the fully opened state, normalized with its range of motion $d_{goal}=\frac{||j_{end}-j_{closed}||}{||j_{opened}-j_{closed}||}$ where $j_{end}$ is the joint angle at the end of the policy rollout, $j_{closed}$ and $j_{opened}$ are the fully closed angle and the fully opened angle respectively. We evaluate each test object for 5 times.
@@
Table 1: Success Rate Metric Results ($\uparrow$): Amount of success trials (normalized distance $<$ 0.1) of different objects’ categories after a full rollout across different methods. The higher the better. $\textbf{AVG}_{c}$ and $\textbf{AVG}_{s}$ refers to metric averaged over categories or samples. For the policy column, $SG$ refers to switch grasp point, $CC$ refers to consistency check and $HF$ refers to history filter. We only add history filter in history-aware methods, and add consistency check with diffusion-based methods because inconsistent predictions in deterministic methods can’t be refined even with re-predictions.
@@
Comparing the $Baselines$ models with $Ours$ in Tables 1 and 2 (see Appendix for normalized distance metrics), we can see that our full method FlowBotHD with the complete set of policy improvements (switch grasp point + consistency check + history filter) achieves the best overall performance. Specifically, compared to the original FlowBot3D [1] model, it increases the success rate averaged across all samples from 87% to 95%, or a 61% reduction in error rate (from 13% to 5%). We see an especially large increase in success rate in the door category (from 23% to 77%), which is the most ambiguous category when fully closed. These quantitative results demonstrate our model’s ability to overall improve the performance and more importantly, better handle multi-modal cases.
Figure 2: Occlusion analysis: We quantitatively demonstrate our model’s performance improvements over severely occluded case. We open the object to different angles, make predictions and plot the RMSE metric ($\downarrow$) against the open ratio. We also include flow visualizations to intuitively show the quality comparison of the predictions.
@@ 6.2 Ablations
We carried out ablations in which we remove various components of the model: No history, no diffusion, and varying whether we use switch grasping (“SG” column), consistency check (“CC” column), and history filtering (“HF” column). We observe large performance improvements with adding history awareness to the diffusion model (comparing Ours to “Ours- No History”). This improvement demonstrates that the extra guidance of history is beneficial for diffusion to reduce the multi-modality in cases of past success in which the opening direction is now disambiguated.
These tables show that our policy improvements of switch grasp point, consistency check, and history filter all contribute to the model performance. The importance of the consistency check potentially means that one challenge for using the current diffusion method is its randomness and, as a result, inconsistency. By heuristically filtering out inconsistent predictions, our model can produce an improved quality of predictions.
@@ 6.3 Qualitative Analysis
To understand the ability of our model to handle occlusions which can not be intuitively shown in the quantitative results, in Figure 2, we show the model predictions for different opening angles. At certain angles, the observed point cloud is severely occluded to become almost a line. We can see from the RMSE plot that the FlowBot3D prediction suddenly fails at a certain angles for which there are severe occlusions. The FlowBotHD distribution looks much better where it still has high probability of predicting the correct direction even at the point of occlusion. The flow visualization shows a clear improvement between our model and the baseline for this ambiguous case.
@@ 6.4 Real-World Experiments
We performed real-world experiments on general articulated objects and specifically our custom made multi-modal door to demonstrate FlowBotHD’s ability to function in real-world scenarios. For the multi-modal door, we demonstrate our model’s ability to open the door in all 4 possible modes, and also the ability to adjust to better grasp points with the switch grasp point policy. Static visualizations of the real-world roll-outs on custom made door and general articulated objects are presented in Figure 3 and Appendix Figure 14, and full videos can be found on our project website. Details of the environmental set-ups including the design of custom made door can be found in Appendix F.
Figure 3: Demonstration of FlowBotHD opening custom made door in different configurations with or without the need to switch grasp point. The first two rows show examples of our model’s successful trials with the first prediction aligned with the current door configuration. The last row shows switch grasp point when the first prediction falls into misaligned modes.
@@ 7 Conclusion
In this work, we propose a history-aware diffusion model FlowBotHD to solve ambiguities in articulated object manipulation, along with a carefully designed heuristic manipulation policy. Both quantitative experiments and qualitative analysis have shown the overall effectiveness of our model and specific improvements over multi-modal or severely occluded cases.
Limitations: While our model greatly improves the ability of the policy to handle ambiguities, we have observed trade-offs in accuracy and consistency when adopting a diffusion pipeline. Our diffusion model still sometimes preserve multi-modality even when the object is opened enough to indicate a deterministic opening direction. While we currently solve this issue by adding a consistency check in the policy, an ideal model should learn to disambiguate by itself based on history trials. Therefore, we hope to further explore how to exert better control over the diffusion’s randomness for better consistency and accuracy.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under Grant No.IIS-2046491 and Toyota Research Institute.
@@ Appendix A Training Details
Model Structure: FlowBotHD consists of a current state encoder, a history state encoder and a denoiser. Both current and history state encoder are based on PointNet++ [39], where the history encoder only includes the PointNet++ encoding module (i.e. no decoder) that outputs a global latent of 128 dimensions. We kept the same model architecture and hyperparameter set as the original PointNet++ paper except for using a regression head for the history encoder output instead of the original segmentation head. We inject the history latent into every layer of the PointNet++ framework in the decoder (see Figure 1). This is performed through a hadamard product of the current point cloud encoding with the latent encodings at every decoding layer to produce the final output of the model. We base the denoiser on a DiT with 5 layers, 4 heads, and a hidden size of 128.
@@
Training Process: We train the model for 450 epochs using AdamW optimizer with a weight decay of 1e-5. We use a learning rate of 1e-4, and a batch size of 128. We evaluate the model every 20 epochs, and save the checkpoint with the best winner-take-all RMSE metric. The Winner-take-all metric means that we repeatedly make predictions for each sample 20 times and record the best RMSE.
@@ Appendix B Simulation Details
We implement a simulation environment with PyBullet that simulates operating a suction gripper to operate PartNet-Mobility objects. We first roll out the simulation with ground truth flow and filter out samples that can’t open with the ground truth motion, due to errors in the simulator. We repeat each sample for 5 times during evaluation. The final test object categories we evaluated on after filtering and repeating are listed in Table 2.
@@
To manipulate the objects, we first attach the suction gripper to the object. This is implemented by first teleporting the gripper to a position close to the target grasp point and then gradually moving towards the target grasp point until contact is detected. After contact, we apply a physical force constraint between the gripper and the object. The move action is then implemented as applying a certain velocity along the predicted direction.
Hyperparameter in Policy: There are 3 hyperparameters to set in Algorithm 1: the threshold for switch grasp point $\delta l$, the threshold for consisteny check $\delta\theta_{cc}$ and the threshold for good movement $\delta m$. We set $\delta l=0.2$, $\delta\theta_{cc}=30^{\circ}$ and $\delta m=$1e-2 in simulation. Table 3 demonstrates the quantitative simulation results for the normalized distance metric, computed in the same way as in Eisner et al. [1].
@@
We demonstrate the simulation process in Fig 4. We visualize the flow predictions (the first row), the simulation process (the last row, x-axis is the step number, the y-axis is the open ratio), and the trajectory plot along with policy signals (the middle plot). In these visuals, we can see our model and policy’s pattern of opening an object: make trials at the beginning, once the door is opened a bit, make consistent predictions based on past history and consistency check.
The first example is quite smooth, our model succeeds to move the door at the first step. Then it continues to make consistent and history-aware predictions until the door is fully opened. Consistency check didn’t have to filter out much predictions due to the good prediction quality, and history is always updated at each step, meaning that we’re always using the previous step’s history for the current step’s prediction. In the second example, our model makes several trials at the fully closed state. History is not updated during these steps and remains none. Once the door is opened, the history information and the consistency check enable the model to execute consistent actions. We can see from the background bar plot that for step 11, the consistency check filters out 12 inconsistent predictions, demonstrating the effectiveness of applying consistency check. Also, we can see that at step 6, the door is opened a bit, but the history is not updated due to the small movement (indicating the prediction is not good enough) which demonstrates how the history filter works.

### v2

@@ 4.3 Articulation Manipulation Policy
Along with the history-aware diffusion model, we propose a novel articulation manipulation policy that utilizes the flow predicting model in real manipulation tasks, taking into consideration the multi-modality of diffusion and the incorporation of object history.
@@
Switch Grasp Point (SG): Following Eisner et al. [1], to achieve the best manipulation efficiency with a suction gripper, we want to grasp at the point with the greatest leverage; thus the optimal action given the predicted flow is to grasp the point that has the maximum flow norm [1]. Due to the multi-modality of diffusion, the prediction at different steps will vary; each prediction might imply a different point that the robot should grasp for maximum leverage. In order to correct for past potential grasping mistakes as well as to maintain a reasonable re-grasping cost, we propose to switch the grasp point whenever the leverage (defined as the current largest flow norm) of the newly predicted grasp point is larger than that of the current grasp point by a threshold: $||\hat{\mathbf{F}}_{max}||-||\hat{\mathbf{F}}_{g}||>\epsilon_{l}$, where $\hat{\mathbf{F}}$ is the current step’s flow prediction for all points in the point cloud, $g$ is the index of the point that is closest to the current grasp point, and $\epsilon_{l}$ is the threshold which we set to 0.2 in practice.
@@
Consistency Check (CC): Learned diffusion models may not always perfectly represent the true distribution over history-conditioned 3DAF values; for example, they might predict multiple modes even when there is unambiguously only a single mode. To handle this, we perform a consistency check for every prediction after the first. Assuming that our model will sample the true mode given enough samples for any observation, we design a heuristic consistency check to filter the actions that are not consistent with the previously executed motion. Specifically, we define the consistency direction $g_{\text{his}}$ as the measured gripper motion caused by executing previous selected flow $\hat{f}_{\text{his}}$ (we only keep the previous execution if its motion was greater than 00). For a subsequent step to be considered “consistent”, the predicted action’s direction shouldn’t deviate too much from that of the previous step; thus we set an angle threshold $\epsilon_{\theta}$ for the consistency check: $\arccos{\frac{\hat{f}_{\text{max}}\cdot g_{\text{his}}}{||\hat{f}_{\text{max}}||\cdot||g_{\text{his}}||}}<\epsilon_{\theta}$, where $\hat{f}_{\text{max}}$ is the current maximum sampled flow direction $\hat{f}_{\text{max}}=\max_{i}\hat{f}_{i}$. We sample 3DAF flows from the diffusion model and compute maximum-leverage actions [1] until we find one that satisfies the consistency check, before executing the sampled action.
@@
History Filter (HF): If we rely solely on a consistency check for next actions, the model may end up in a degenerate state where its previous predictions are incorrect, enforcing that only incorrect predictions are permitted by consistency. Therefore, we include a history filter that only updates the history point cloud $\textbf{P}_{\text{his}}$ and history flow $\mathbf{F}_{\text{his}}$ if the previous action successfully achieves reasonable gripper motion without losing contact. Specifically, after an execution, we only update the history if two criteria are satisfied:
@@
    The measured gripper motion $g_{\text{cur}}$ caused by the currently-executed action is consistent with past executions: $\arccos{\frac{g_{\text{cur}}\cdot g_{\text{his}}}{||g_{\text{cur}}||\cdot||g_{\text{his}}||}}<\epsilon_{c}=90^{\circ}$ where $g_{\text{cur}}$ is the current gripper movement and $g_{\text{his}}$ is the gripper movement from the previous execution.
@@
    The gripper motion $g_{\text{cur}}$ is larger than a threshold $||g_{\text{cur}}||>\epsilon_{m}$ where $g_{\text{cur}}$ is the current gripper movement and $\epsilon_{m}$ is the threshold
@@
     $\mathbf{P}_{\text{obs}}\leftarrow$ Observation
     $\hat{\mathbf{F}}\sim D_{\theta}(\mathbf{P}_{\text{obs}},\mathbf{S}_{\text{his}})$   // Sample the current flow direction from the diffusion model
     $\hat{f}_{\text{max}}\leftarrow\texttt{GetMaxFlow}(\hat{\mathbf{F}})$    // Find the point with the max predicted flow [1]
@@
     $p_{\text{grasp}}\leftarrow\texttt{SelectGraspPoint}(\hat{\mathbf{F}})$
@@
        $p_{\text{grasp,his}}=p_{\text{grasp}}$, ;
@@
        $\texttt{Grasp}(p_{\text{grasp}})$ ;
@@
     $g_{\text{cur}}\leftarrow\texttt{GetDeltaGripper}()$
     if $\texttt{Consistent}(g_{\text{cur}},g_{\text{his}})$  then
        $g_{\text{his}}\leftarrow g_{\text{cur}}$
        if $||g_{\text{cur}}||>\epsilon_{m}$ then
           $\mathbf{S}_{\text{his}}\leftarrow(\mathbf{P}_{\text{obs}},\hat{\mathbf{F}})$ ;
@@ 5.1 Simulation results
To evaluate each model’s ability to open objects, we construct a simulation environment in the PyBullet [42] simulator consisting of a suction gripper and objects from the PartNet-Mobility dataset, where the environment’s goal is to successfully articulate the object. We initialize every object as fully closed and filter out instances that cannot open from the fully closed state even with the ground truth action sequence (e.g. due to self-collision). Following prior work [1], we compute 2 evaluation metrics: Success Rate: the percent of objects the model succeeds to open $>90\%$ of the joint’s full range of motion, and Normalized Distance [13]: the mean distance between the end state of this trial and the fully opened state, normalized with its range of motion $d_{\text{goal}}=\frac{||j_{\text{end}}-j_{\text{closed}}||}{||j_{\text{opened}}-j_{\text{closed}}||}$ where $j_{\text{end}}$ is the joint angle at the end of the policy rollout, $j_{\text{closed}}$ and $j_{\text{opened}}$ are the fully closed angle and the fully opened angle respectively. We evaluate each test object for 5 trials.
@@
Table 1: Success Rate Metric Results ($\uparrow$): Fraction of successful trials (normalized distance $<$ 0.1) of different object categories after a full rollout across different methods; the higher the better. $\textbf{AVG}_{c}$ and $\textbf{AVG}_{s}$ refers to metric averaged over categories or samples. For the policy column, $SG$ refers to switch grasp point, $CC$ refers to consistency check and $HF$ refers to history filter. We only add history filter in history-aware methods and only add consistency check with diffusion-based methods because inconsistent predictions in deterministic methods can’t be refined even with re-predictions.
@@
Comparing the Baseline models with Ours in Tables 1 and 3 (see Appendix for normalized distance metrics), we can see that our full method (FlowBotHD) with the complete set of policy improvements (switch grasp point + consistency check + history filter) achieves the best overall performance. Specifically, compared to the original FlowBot3D [1] model, it increases the success rate averaged across all samples from 87% to 95%, or a 61% reduction in error rate (from 13% to 5%). We observe a particular increase in success rate in the door category (from 23% to 77%), which contains the most visual ambiguities of any category when instances are fully closed. These quantitative results demonstrate our model’s ability to overall improve the performance and more importantly, to better handle multi-modal cases.
Figure 3: We demonstrate our model’s performance improvement over the baseline in a severely occluded case. We open the object to different angles, make predictions, and plot the RMSE metric ($\downarrow$) against the open ratio. The baseline (top) fails at the point of severe occlusion, whereas our method (bottom) continues to make stable predictions using history. We also include flow visualizations (right) to show the quality of the predictions in both cases.
@@ 5.2 Ablations
To understand the impact of our design decisions, we perfomed ablation studies in which we remove various components of the model: No history, no diffusion, and varying whether we use switch grasping (“SG” column), consistency check (“CC” column), and history filtering (“HF” column). Looking at Tables 1 and 3, we observe large performance improvements when including history awareness in the diffusion model (comparing Ours to “Ours- No History”). This improvement demonstrates that the extra guidance of history is beneficial for diffusion to reduce the multi-modality in cases of past success in which the opening direction is now disambiguated.
These tables show that our policy improvements of switching grasp point, applying a consistency check, and applying a history filter all contribute to increased model performance. The importance of the consistency check potentially implies that one challenge for using diffusion models for visual manipulation policies is randomness and, as a result, inconsistency. By filtering out inconsistent predictions, our policy takes actions that lead to higher success rates than without filtering.
@@ 5.3 Qualitative Analysis
To understand the ability of our model to handle occlusions which are not immediately apparent from our quantitative metrics, in Figure 3, we visualize the model predictions for different door opening angles. At certain angles, the point cloud is severely occluded such that only the vertical edge of the door is visible from the camera’s perspective. We can see from the RMSE plot that FlowBot3D consistently mispredicts the 3DAF on examples where the opening angle causes severe occlusions. FlowBotHD performs substantially better in these occluded cases by leveraging history, predicting the correct direction even at the point of occlusion. The flow visualization shows a clear improvement between our model and the baseline for this ambiguous case.
@@ 5.4 Real-World Experiments
We performed real-world experiments on general articulated objects, as well as a custom made multi-modal door to demonstrate FlowBotHD’s ability to function in real-world ambiguous scenarios (Figures 1 and 4). For the multi-modal door, we demonstrate our model’s ability to open the door in all 4 possible modes, and also the ability to adjust to better grasp points with the switch grasp point policy. Static visualizations of the real-world roll-outs on our custom-made door and general articulated objects are presented in Figure 4 and Appendix Figure 15, and full videos can be found on our project website. Details of our real-world setup can be found in Appendix F.
Figure 4: Demonstration of FlowBotHD opening a custom made door in different configurations. The first two rows show examples in which the model’s initial predictions aligned with the current door configuration. The last row shows an example in which the model’s initial prediction fails and the model switches its grasp point, ultimately leading to a success.
@@ 6 Conclusion
In this work, we propose a history-aware diffusion model FlowBotHD to handle ambiguities encountered while manipulating articulated objects, along with a carefully designed heuristic manipulation policy. Both quantitative experiments and qualitative analysis show the overall effectiveness of our model and specific improvements in multi-modal or severely occluded cases.
Limitations: While our model greatly improves the ability of the policy to handle ambiguities, we have observed trade-offs in accuracy and consistency when building on top of a diffusion pipeline. Our diffusion model sometimes makes multimodal predictions even when the object is opened enough to disambiguate the true opening direction. While we currently mitigate this issue by adding a consistency check in the policy, an ideal model would learn to disambiguate perfectly based solely history. Therefore, we hope to further explore how to exert better control over the diffusion’s learned distributions for better consistency and accuracy.
@@ Acknowledgments
This material is based upon work supported by the National Science Foundation under NSF CAREER Grant No. IIS-2046491 and Toyota Research Institute.
@@ Appendix A Training Details
Model Structure: FlowBotHD consists of a current state encoder $g_{\phi}$, a history state encoder $h_{\gamma}$ and a denoiser $\mathbf{D}_{\theta}$. Both current and history state encoder are based on PointNet++ [39], where the history encoder only includes the PointNet++ global encoding module that outputs a global latent of 128 dimensions. We kept the same model architecture and hyperparameter set as the original PointNet++ paper. We inject the history latent into every layer of the PointNet++ framework in the decoder (see Figure 2). This is performed through a Hadamard product of the current point cloud encoding with the latent encodings at every decoding layer to produce the final output of the model. We base the denoiser on a DiT with 5 layers, 4 heads, and a hidden size of 128.
@@
Training Process: We set diffusion timestep to be 100 steps. We train the model for 450 epochs using AdamW optimizer with a weight decay of 1e-5. We use a learning rate of 1e-4, and a batch size of 128. We evaluate the model every 20 epochs, and save the checkpoint with the best winner-take-all RMSE metric. The Winner-take-all metric means that we repeatedly make predictions for each sample 20 times and record the best RMSE.
@@ Appendix B Simulation Details
We design a simulation environment in the PyBullet simulator that simulates operating a suction gripper to operate PartNet-Mobility objects. We first roll out the simulation with ground truth flow and filter out samples that can’t open with the ground truth motion, due to errors in the simulator. We repeat each sample for 5 times during evaluation. The final test object categories we evaluated on after filtering and repeating are listed in Table 2.
@@
To manipulate the objects, we first attach the suction gripper to the object. This is achieved by first teleporting the gripper to a position close to the target grasp point and then gradually moving towards the target grasp point until contact is detected. After contact, we apply a physical force constraint between the gripper and the object. The move action is then implemented as applying a certain velocity along the predicted direction.
Hyperparameter in Policy: There are 3 hyperparameters to set in Algorithm 1: the threshold for switch grasp point $\epsilon_{l}$, the threshold for consisteny check $\epsilon_{\theta}$ and the threshold for good movement $\epsilon_{m}$. We set $\epsilon_{l}=0.2$, $\epsilon_{\theta}=30^{\circ}$ and $\epsilon_{m}=$1e-2 in simulation. Table 3 demonstrates the quantitative simulation results for the normalized distance metric, computed in the same way as in Eisner et al. [1].
@@
We demonstrate the simulation process in Fig 5. We visualize the flow predictions (the first row), the simulation process (the last row, x-axis is the step number, the y-axis is the open ratio), and the trajectory plot along with policy signals (the middle plot). In these visuals, we can see our model and policy’s pattern of opening an object: make one or more attempts at opening at the beginning of the episode, and once the door is opened a bit, make consistent predictions based on past history and consistency check.
The first example is quite smooth: our model succeeds in moving the door on the first step. It continues to make consistent and history-aware predictions until the door is fully opened. Qualitatively, the consistency check didn’t filter out many predictions due to high prediction quality, and due to the smooth execution the history is updated at each step, meaning that the policy always uesd the previous step’s history for the next step’s prediction. In the second example, our model makes several attempts to open the object when in the fully closed state. The observation history is not updated during these steps. Once the door is opened a bit, the history information and the consistency check enable the model to execute consistent actions. In the bar plot, we se that for step 11, the consistency check filters out 12 inconsistent predictions, demonstrating the effectiveness of applying a consistency check. Also, we can see that at step 6, the door is opened a bit, but the history is not updated due to the small movement (indicating the prediction is not good enough) which demonstrates the importance of tuning the history filter.